# A Fast and Model Based Approach for Evaluating Task-Competence of Antagonistic Continuum Arms

### v3

@@ I INTRODUCTION
Fluid-driven soft robot arms seek to capture the physical intelligence of muscular hydrostats, such as elephant trunks and octopus arms [1], in order to improve robot robustness and safety around humans [2]. These arms combine soft pneumatic actuators in parallel and in series to produce robotic arms that can bend in any direction at multiple points (see example in Fig. 1) [3, 4, 5, 6]. Recent works have taken major steps toward soft robotic arms that can complete human-scale tasks, exploring backbone-free antagonistic designs for stiffness control [7, 8, 9] and demonstrating contact rich tasks such as opening drawers [6], washing human subjects [10] and assisted eating [11]. However, arm demonstrations are consistently limited to small external loads, even with high actuator pressures [3, 8, 12, 6, 11].
The mechanical reason for limited force is not obvious, but it is evident even in recent, novel demonstrations: in Jiang et al., no tasks are demonstrated with additional weights [6], and in related work the arm’s maximum characterized load is 2.8N [6][12]. The arm in Guan et al. is shown, in a separate work, to deform its entire body length under 4N of tip loading [11, 13]. Investigations of the mapping between design and performance, e.g., load capacity across the workspace, are hampered by the lack of model-based design tools for soft robotic arm. Soft arm modeling efforts have focused on control, and use experimental, homogenized stiffness parameters that are specific to the design being tested [14, 15, 16]. Practically, an arm’s ability to complete a desired task has been determined by building and testing it.
@@
Figure 2: Overview of our mechanics model. (A): A soft robot arm composed of two bellows and two muscles can be fully parameterized by its centerline twists $\accentset{\scriptscriptstyle{\circleright}}{g}_{oi}$ and transformations between actuators ${}_{o}g_{o\alpha}$. (B): When a tip load $\mathbf{q}_{\textrm{tip}}$ is applied, wrenches $\mathbf{q}_{i}$ are induced along the arm’s backbone. To achieve static equilibrium, each actuator contributes a reaction force and moment to balance the load. (C): We consider two types of actuators: contracting McKibben artificial muscles and extensile bellows actuators. The characterized force functions $f(\epsilon,p)$ of each actuator are also shown, with actuation regimes labeled according to [17].
Model-based design tools require model formulations that generalize across designs and methods of using those models to produce informative, interpretable results. Prior work [17] has developed generalizable models of soft arms, but significant gaps remain in developing approaches that utilize these models to provide insights about the mapping between arm design and task-specific performance.
@@ II CONTINUUM MODEL
This section briefly reviews our underlying assumptions, notation, and mechanics-based continuum model, which is a planar Cosserat-rod formulation of the design-oriented model of [17]. Our work makes the following assumptions:
@@
    Design: We assume the arm and actuators act as quasi-static shear-free, torsion-free rods. Prior work has shown that while this may not be valid in minimal cases (three long, thin actuators), it becomes reasonable when stiffer, shorter or more actuators are used [18].
@@ II-A Cosserat rod kinematics
Using eqn. 3, we can describe the poses along each actuator ${}_{\alpha}g=\{{}_{\alpha}g_{1},{}_{\alpha}g_{2},\cdots{}_{\alpha}g_{N+1}\}$ using the base pose ${}_{\alpha}g_{1}$ and the twists of each segment ${}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}=\{{}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}_{1},{}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}_{2},\cdots,{}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}_{N}\}$, which are more closely related to the mechanics (Fig. 2A). Furthermore, our assumption of a rigid and constant cross-section enables us relate the twists of each actuator to the twists of the manipulator centerline. Let the pose and twists of the manipulator centerline at node $i$ be ${}_{o}g_{i}$ and ${}_{o}\accentset{\scriptscriptstyle{\circleright}}{g}_{i}$ respectively, and let ${}_{o}g_{o\alpha}$ be the transformation from centerline $o$ to actuator $\alpha$ in each cross-section, which is usually $\begin{bmatrix}0&r_{\alpha}&0\end{bmatrix}^{\textrm{T}}$. As shown in [19], the twist of any actuator can be computed from the centerline twist as:
@@ II-C Internal Reaction
$\mathbf{a}_{i}=\begin{bmatrix}1&1&\dots&1\\ 0&0&\dots&0\\ r_{1}&r_{2}&\dots&r_{m}\end{bmatrix}\begin{bmatrix}f_{A}({}_{A}\epsilon_{i},p_{A})\\ f_{B}({}_{B}\epsilon_{i},p_{B})\\ \vdots\\ f_{\mu}({}_{\mu}\epsilon_{i},p_{\mu})\end{bmatrix}+\\ 1\textrm{e}5\begin{bmatrix}{}_{A}\gamma_{i}\\ {}_{B}\gamma_{i}\\ \vdots\\ {}_{\mu}\gamma_{i}\end{bmatrix}+\begin{bmatrix}0&0&\dots&0\\ 0&0&\dots&0\\ 1&1&\dots&1\end{bmatrix}\begin{bmatrix}\tau_{A}(\kappa_{A},p_{A})\\ \tau_{B}(\kappa_{B},p_{B})\\ \vdots\\ \tau_{\mu}(\kappa_{\mu},p_{\mu})\end{bmatrix}$ (7)
@@ III-C1 Absolute Attainability
These propositions are consequences of our actuator models being continuous and monotonic, which preserve topological invariants like boundaries and connectedness [20] (see SI for proof). We can therefore approximate whether $\mathbf{w}^{*}_{i}$ is in $\mathbf{a}_{i}[\mathcal{P}]$ by sampling a set of points $\mathbf{E}$ from the edges of $\partial\mathcal{P}$ and checking if $\mathbf{w}^{*}_{i}$ is enclosed by the convex hull $\textrm{Conv}(\mathbf{a}_{i}[{\mathbf{E}}])$. We call this $\mathcal{H}_{i}$, the $\mathit{\mathbf{i}}$-th absolute wrench hull, which also visualizes the possible reactions across $\mathcal{P}$. Across all $i$, we will refer to the collection $\mathcal{H}=(\mathcal{H}_{1},\dots,\mathcal{H}_{N})$ as the absolute wrench hull, and if each $\mathbf{w}_{i}^{*}$ is enclosed in each $\mathcal{H}_{i}$ then we say the task $({}_{o}\accentset{\scriptscriptstyle{\circleright}}{g},\mathbf{q}_{\textrm{tip}}^{*})$ is absolutely attainable (Fig. 3D). If any $\mathbf{w}^{*}_{i}$ are not in $\mathcal{H}_{i}$, the shortest distance between them can be found by solving a convex quadratic program described in our SI, and we will refer to their sum as the absolute unattainability.
@@ III-C2 Relative Attainability
Figure 5: Attainability methods accurately predict the ability of arm designs to accomplish a range of shared tasks. Reaction wrench requirements are imposed by 67 sampled tip-loads applied to two arm designs held at a task shape. The attainable wrench hull of an antagonistic arm encloses significantly more wrench requirements than that of an arm with only bellows, and this is quantified in (B) and (C) - vertical lines are medians. Attainability thus predicts that the antagonistic arm can complete more of the tasks, and this is confirmed by exhaustively searching in (D) and (E) for each arm’s best attempt to match the specified shape while subject to each of the sampled tip loads . Tip loads are plotted without moments, and were uniformly sampled from the range of [±$\begin{bmatrix}\pm 10\textrm{N}&\pm 10\textrm{N}&\pm 1\textrm{Nm}\end{bmatrix}$.
@@ IV-A Experimental Characterization of Actuators
The proposed arm designs analyzed below are actuated by either McKibben actuators, bellows actuators or both. The force function $f(\epsilon,p)$ of the McKibben actuator is taken from [17], while the force function of the bellows actuator is measured via a similar process, but is novel to this work - see our SI for details. The bending moment function $\tau(\kappa,p)$ is adapted from findings in [21] - we assume a linear stiffness that varies with pressure, and use the following function:
@@
While actuator force models were extracted from available data, bending stiffness parameters were estimated due to lack of suitable test data. We selected $K=-0.285\ \mathrm{Nm^{2}}$ for both McKibbens and bellows, and use $\bar{p}=50\mathrm{kPa}$ for bellows and $100\mathrm{kPa}$ for McKibbens. The presented method is valid for any bending stiffness model that meets the listed assumptions, and further work is needed to identify a robust, generalizable bending stiffness model for soft arms.
@@ IV-C Evaluation of Antagonistic Actuation
The antagonistic arm has the lowest absolute unattainability for the first two task-shapes, followed by the bellows-only arm, and finally the McKibben-only arm. Comparing the absolute wrench-hulls of each arm suggests why the antagonistic arm is more capable of the two: because the antagonistic arm combines extending and contracting actuators, its absolute wrench-hull is closer to the origin, and therefore better at enclosing requirement wrench sequences, which are centered around the origin. The bellows-only arm may have a larger wrench-hull, but because it only has extending actuators, its wrench-hull only extends to the positive reaction force half-plane. Likewise the McKibben-only arm, whose actuators can only contract, has a wrench-hull entirely with negative reaction-force. Passive actuator reaction forces do provide countering force, as noted in [18], but are strain-dependent and small in magnitude, while antagonistic countering forces are controllable via pressure and can be large in magnitude.

### v4

@@ I INTRODUCTION
Fluid-driven soft robot arms seek to capture the physical intelligence of muscular hydrostats, such as elephant trunks and octopus arms [kier_1985], in order to improve robot robustness and safety around humans [chen_review_2022]. These arms combine soft pneumatic actuators in parallel and in series to produce robotic arms that can bend in any direction at multiple points (see example in Fig. 1) [octarm, davis_bidirectional_varstiff, stiff_flop_jamming, jiang_hierarchical_2021]. Recent works have taken major steps toward soft robotic arms that can complete human-scale tasks, exploring backbone-free antagonistic designs for stiffness control [var_grip_stiffness_independence, davis_varstiff_advancedrobotics, OlsonAsselmeier_RoboSoft] and demonstrating contact rich tasks such as opening drawers [jiang_hierarchical_2021], washing human subjects [zlatintsi_i-support_2020] and assisted eating [guan_trimmed_2023]. However, arm demonstrations are consistently limited to small external loads, even with high actuator pressures [octarm, davis_varstiff_advancedrobotics, jiang_design_2016, jiang_hierarchical_2021, guan_trimmed_2023].
The mechanical reason for limited force is not obvious, but it is evident even in recent, novel demonstrations: in Jiang et al., no tasks are demonstrated with additional weights [jiang_hierarchical_2021], and in related work the arm’s maximum characterized load is 2.8N [jiang_hierarchical_2021][jiang_design_2016]. The arm in Guan et al. is shown, in a separate work, to deform its entire body length under 4N of tip loading [guan_trimmed_2023, stella_piecewise_2023_s]. Investigations of the mapping between design and performance, e.g., load capacity across the workspace, are hampered by the lack of model-based design tools for soft robotic arm. Soft arm modeling efforts have focused on control, and use experimental, homogenized stiffness parameters that are specific to the design being tested [renda_cosserat_tro, trivedi_model_tro, rus_ijrr_control]. Practically, an arm’s ability to complete a desired task has been determined by building and testing it.
@@
Figure 2: Overview of our mechanics model. (A): A soft robot arm composed of two bellows and two muscles can be fully parameterized by its centerline twists $\accentset{\scriptscriptstyle{\circleright}}{g}_{oi}$ and transformations between actuators $_{o}g_{o\alpha}$. (B): When a tip load $\mathbf{q}_{\textrm{tip}}$ is applied, wrenches $\mathbf{q}_{i}$ are induced along the arm’s backbone. To achieve static equilibrium, each actuator contributes a reaction force and moment to balance the load. (C): We consider two types of actuators: contracting McKibben artificial muscles and extensile bellows actuators. The characterized force functions $f(\epsilon,p)$ of each actuator are also shown, with actuation regimes labeled according to [olson_eulerbernoulli_2020].
Model-based design tools require model formulations that generalize across designs and methods of using those models to produce informative, interpretable results. Prior work [olson_eulerbernoulli_2020] has developed generalizable models of soft arms, but significant gaps remain in developing approaches that utilize these models to provide insights about the mapping between arm design and task-specific performance.
@@ II CONTINUUM MODEL
This section briefly reviews our underlying assumptions, notation, and mechanics-based continuum model, which is a planar Cosserat-rod formulation of the design-oriented model of [olson_eulerbernoulli_2020]. Our work makes the following assumptions:
@@
    Design: We assume the arm and actuators act as quasi-static shear-free, torsion-free rods. Prior work has shown that while this may not be valid in minimal cases (three long, thin actuators), it becomes reasonable when stiffer, shorter or more actuators are used [me_mocap].
@@ II-A Cosserat rod kinematics
Using eqn. 3, we can describe the poses along each actuator $_{\alpha}g=\{{}_{\alpha}g_{1},{}_{\alpha}g_{2},\cdots{}_{\alpha}g_{N+1}\}$ using the base pose $_{\alpha}g_{1}$ and the twists of each segment $_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}=\{{}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}_{1},{}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}_{2},\cdots,{}_{\alpha}\accentset{\scriptscriptstyle{\circleright}}{g}_{N}\}$, which are more closely related to the mechanics (Fig. 2A). Furthermore, our assumption of a rigid and constant cross-section enables us relate the twists of each actuator to the twists of the manipulator centerline. Let the pose and twists of the manipulator centerline at node $i$ be $_{o}g_{i}$ and $_{o}\accentset{\scriptscriptstyle{\circleright}}{g}_{i}$ respectively, and let $_{o}g_{o\alpha}$ be the transformation from centerline $o$ to actuator $\alpha$ in each cross-section, which is usually $\begin{bmatrix}0&r_{\alpha}&0\end{bmatrix}^{\textrm{T}}$. As shown in [fan_linear_2023], the twist of any actuator can be computed from the centerline twist as:
@@ II-C Internal Reaction
$\mathbf{a}_{i}=\;\begin{bmatrix}1&1&\dots&1\\ 0&0&\dots&0\\ r_{1}&r_{2}&\dots&r_{m}\end{bmatrix}\begin{bmatrix}f_{A}({}_{A}\epsilon_{i},p_{A})\\ f_{B}({}_{B}\epsilon_{i},p_{B})\\ \vdots\\ f_{\mu}({}_{\mu}\epsilon_{i},p_{\mu})\end{bmatrix}$+ (7)
$1e3\begin{bmatrix}0&0&\dots&0\\ 1&1&\dots&1\\ 0&0&\dots&0\end{bmatrix}\begin{bmatrix}{}_{A}\gamma_{i}\\ {}_{B}\gamma_{i}\\ \vdots\\ {}_{\mu}\gamma_{i}\end{bmatrix}+\begin{bmatrix}0&0&\dots&0\\ 0&0&\dots&0\\ 1&1&\dots&1\end{bmatrix}\begin{bmatrix}\tau_{A}({}_{A}\kappa_{i},p_{A})\\ \tau_{B}({}_{B}\kappa_{i},p_{B})\\ \vdots\\ \tau_{\mu}({}_{\mu}\kappa_{i},p_{\mu})\end{bmatrix}$
@@ III-C1 Absolute Attainability
These propositions are consequences of our actuator models being continuous and monotonic, which preserve topological invariants like boundaries and connectedness [munkres_topology_2000] (see SI for proof). We can therefore approximate whether $\mathbf{w}^{*}_{i}$ is in $\mathbf{a}_{i}[\mathcal{P}]$ by sampling a set of points $\mathbf{E}$ from the edges of $\partial\mathcal{P}$ and checking if $\mathbf{w}^{*}_{i}$ is enclosed by the convex hull $\textrm{Conv}(\mathbf{a}_{i}[{\mathbf{E}}])$. We call this $\mathcal{H}_{i}$, the $\mathit{\mathbf{i}}$-th absolute wrench hull, which also visualizes the possible reactions across $\mathcal{P}$. Across all $i$, we will refer to the collection $\mathcal{H}=(\mathcal{H}_{1},\dots,\mathcal{H}_{N})$ as the absolute wrench hull, and if each $\mathbf{w}_{i}^{*}$ is enclosed in each $\mathcal{H}_{i}$ then we say the task $({}_{o}\accentset{\scriptscriptstyle{\circleright}}{g},\mathbf{q}_{\textrm{tip}}^{*})$ is absolutely attainable (Fig. 3D). If any $\mathbf{w}^{*}_{i}$ are not in $\mathcal{H}_{i}$, the shortest distance between them can be found by solving a convex quadratic program described in our SI, and we will refer to their sum as the absolute unattainability.
@@ III-C2 Relative Attainability
Figure 5: Attainability methods accurately predict the ability of arm designs to accomplish a range of shared tasks. Reaction wrench requirements are imposed by 67 sampled tip-loads applied to two arm designs held at a task shape. The attainable wrench hull of an antagonistic arm encloses significantly more wrench requirements than that of an arm with only bellows, and this is quantified in (B) and (C) - vertical lines are medians. Attainability thus predicts that the antagonistic arm can complete more of the tasks, and this is confirmed by exhaustively searching in (D) and (E) for each arm’s best attempt to match the specified shape while subject to each of the sampled tip loads . Tip loads are plotted without moments, and were uniformly sampled from the range of $\begin{bmatrix}\pm 10\textrm{N}&\pm 10\textrm{N}&\pm 1\textrm{Nm}\end{bmatrix}$.
@@ IV-A Experimental Characterization of Actuators
The proposed arm designs analyzed below are actuated by either McKibben actuators, bellows actuators or both. The force function $f(\epsilon,p)$ of the McKibben actuator is taken from [olson_eulerbernoulli_2020], while the force function of the bellows actuator is measured via a similar process, but is novel to this work - see our SI for details. The bending moment function $\tau(\kappa,p)$ is adapted from findings in [garbulinski_bending_2022] - we assume a linear stiffness that varies with pressure, and use the following function:
@@
[…]
@@ IV-C Evaluation of Antagonistic Actuation
The antagonistic arm has the lowest absolute unattainability for the first two task-shapes, followed by the bellows-only arm, and finally the McKibben-only arm. Comparing the absolute wrench-hulls of each arm suggests why the antagonistic arm is more capable of the two: because the antagonistic arm combines extending and contracting actuators, its absolute wrench-hull is closer to the origin, and therefore better at enclosing requirement wrench sequences, which are centered around the origin. The bellows-only arm may have a larger wrench-hull, but because it only has extending actuators, its wrench-hull only extends to the positive reaction force half-plane. Likewise the McKibben-only arm, whose actuators can only contract, has a wrench-hull entirely with negative reaction-force. Passive actuator reaction forces do provide countering force, as noted in [me_mocap], but are strain-dependent and small in magnitude, while antagonistic countering forces are controllable via pressure and can be large in magnitude.